OpenStudy (anonymous):

Can you please show an example of a Polynomial Curve-Fitting with 4 points?

OpenStudy (anonymous):

Any 4 points?

OpenStudy (anonymous):

Any As easy as possible please

OpenStudy (anonymous):

Given the points {a, b, c, d} the polynomial P(x) = (x-a)(x-b)(x-c)(x-d) works.

OpenStudy (anonymous):

How about if you're given (-1,3) (0,5) (1,-1) (2,2)

OpenStudy (anonymous):

Okie dokie, do you know linear algebra/are you comfortable working with matrices?

OpenStudy (anonymous):

Yes, I am. This is our upcoming topic in our class.

OpenStudy (anonymous):

Perfect (otherwise this problem would be a lot harder)! In this case, a degree 3 polynomial will work, \[P(x) = a_0 +a_1x+a_2x^2+a_3x^3\] and our job is to find \[{a_0,a_1,a_2,a_3}\]

OpenStudy (anonymous):

Yes.Then substitue the ff points to the polynomial?

OpenStudy (anonymous):

If you can show the one already at the row echelon form, itcould be great help

OpenStudy (anonymous):

Yup,yup. Since P(x) must pass through each of our points so we obtain the system P( -1 ) = 3 P( 0 ) = 5 P( 1 ) = -1 P( 2 ) = 2

OpenStudy (anonymous):

I see. Please continue, sir.

OpenStudy (anonymous):

K so our equations look like \[a_0 + a_1(-1)+a_2(-1)^2+a_3(-1)^3 = 3\] \[a_0+a_1(0)+a_2(0)^2 + a_3(0)^3 = 5\] \[a_0+a_1(1)+a_2(1)^2+a_3(1)^3 = -1\] \[a_0+a_1(2)+a_2(2)^2+a_3(2)^3 = 2\]

OpenStudy (anonymous):

The matrix will be. 1 -1 1 -1 3 1 0 0 0 5 1 1 1 1 -1 1 2 4 8 2

OpenStudy (anonymous):

In Av=b form, matrix for that would be A v = b |1 -1 -1 -1| |a_0| | 5 | |1 0 0 0| |a_1| = | 3 | |1 1 1 1| |a_2| |-1 | |1 2 2 2| |a_3| | 2 | So we just need to the inverse of A

OpenStudy (anonymous):

Your way was much easier to write it lol.

OpenStudy (anonymous):

i guess so. lol

OpenStudy (anonymous):

I don't know how to work with inverse yet. I'm sorry

OpenStudy (anonymous):

Ahh, well it's the same thing as putting the matrix A into row reduced echelon form

OpenStudy (anonymous):

shouldn't the first row be 1 -1 1 -1 5?

OpenStudy (anonymous):

Oh, I see.

OpenStudy (anonymous):

Whoops, you are correct, I forgot to expand any of the powers, was too focused on making it look like a matrix haha

OpenStudy (anonymous):

Now, we just row reduce this and we're particularly done : D |1 -1 1 -1| 5 | |1 0 0 0 | 3 | |1 1 1 1 | -1 | |1 2 4 8 | 2 |

OpenStudy (anonymous):

a0 = 3.000 a1 = -4.500 a 2 = -1.000 a3 = 1.500

OpenStudy (anonymous):

Is it? :D Thank you for your time sir :D

OpenStudy (anonymous):

Yup yup, just sub those values back in and write your answer as \[P(x) = a_0 +a_1x+a_2x^2+a_3x^3\] and we're done :) Good job!

OpenStudy (anonymous):

Thank you for your help sir :D